What is the rate constant, in s-1, of a first order process that has a half life of 12.1 seconds


1
Expert's answer
2021-07-28T01:33:02-0400


r=k[A]x[B]y

ln2/K = 12.1

K = 0.121/12.1

= 10-2M/s